OpenStudy (anonymous):

Find the angle theta between the vectors. v = 4i - j + k , w= 2i +3j + 5k

OpenStudy (anonymous):

plz do not give me the answer just point me in the right direciton

OpenStudy (anonymous):

Use the property that \(A\cdot B=|A||B|\cos(\theta)\)

OpenStudy (anonymous):

that is exactly what im doing and i keep getting 53.89 degrees which is incorrect.

OpenStudy (anonymous):

I got 10 for the dot product. Is that what y'all got too?

OpenStudy (helder_edwin):

\[ \large v\cdot w=10 \] \[ \large v\cdot v=18\Rightarrow |v|=\sqrt{18} \] \[ \large w\cdot w=38\Rightarrow|w|=\sqrt{38} \] right?

OpenStudy (helder_edwin):

well @m.auld64 ??